I have tried this problem out for the past 30 minutes. It's just that fractions are added in to the linear equation problems and it's confusing me .
Please help me solve this question : (3x +1) /4 - (x+2) / 5 = 2 Thank You

Answer by Alan3354(69443)

(3x +1) /4 - (x+2) / 5 = 2
---------
Multiply by 20
5(3x +1) - 4(x+2) = 40
15x+5 - 4x-8 = 40
11x - 3 = 40
x = 43/11